OpenStudy (anonymous):

25x^2-30x+9 this is undr a square root symbol

OpenStudy (anonymous):

\[\sqrt{?}\]25x^2-30x+9

OpenStudy (saifoo.khan):

\[\sqrt{25x^2-30x+9}\] \[\sqrt{(-5x+3)^2}\] \[|-5x+3|\]